# Richard A Berk

### v2

@@ 7 Applying Genetic Algorithms To The Fitted AIRS Data
A random forest was trained as before using 500 regression trees for the June extreme heat wave. The gain score data used for training were the same as the gain score data employed in the earlier analyses. A trained and averaged ensemble of regression trees can serve as a survival function for a genetic algorithm’s construction of an ideal, synthetic population of grid cells experiencing a true heat wave.
@@
In the same manner, a random forest was trained as before using 500 regression trees on June actual heat wave 1. The gain score data were the same as used in the earlier analyses. A trained and averaged ensemble of regression trees can serve as a different survival function for a genetic algorithm’s construction of an ideal, synthetic population of grid cells experiencing a faux heat wave.
@@
The two survival functions were used as input to the procedure GA in R written by Luca Scrucca (2013). The standard defaults for numeric predictors worked well. For both survival functions, 5000 iterations were specified leading to 5000 synthetic populations with 100 grid cells each having increasing fitness on the average. As a first approximation, one can think of a genetic algorithm as jittering predictor values, computing new fitted values from the survival function using the jittered predictors, saving the subset of the grid cells with the larger gain scores, and repeating this process many times until there is a population that includes only very fit grid cells (i.e., having large gain scores).131313 The GA software in R has several tuning parameters with good default values.(1) For each synthetic population of grid cells, the predictors from the five grid cells with the highest computed surface temperatures were not altered. (2) The crossover probability for each possible pair of predictors was set at 0.8. A crossover refers to a pair of “parent” grid cells combining their predictor values to form an “offspring” grid cell. (3) Because all of the predictors are treated as numeric in this application, the predictor values for an offspring grid cell are computed as the average of the corresponding predictors from a random pair of parent grid cells. (4) The mutation probability for each grid cell’s predictor was set at 0.1. (5) The mutation method used replaced the value of the existing numeric variable with a random value within a specified range drawn at random from a uniform distribution. The size of a synthetic population is not a tuning parameter but 100 grid cells seemed a good compromise between an ability to examine the output visually in detail and having a sufficient number of observations to arrive at sufficiently stable parameter estimates.
@@
The grid cell mean gain score for the final synthetic population produced by the actual heat wave survival function was 7.26 degrees K. The grid cell mean gain score for the final synthetic population produced by the faux heat wave 1 survival function was 5.76 degrees K. Because of differences in the two survival functions, the genetic algorithm could not equate the two. One possible inference, already mentioned, is that the physics for a true heat wave may differ somewhat from the physics of a faux heat wave.141414 Predictor dependence was reduced substantially because of the predictor randomness the genetic algorithm introduced.
@@
Predictor True Heat Wave Faux Heat Wave
@@
Table 2 provides a bit more evidence about the ways in which an ideal true heat wave can differ from an ideal faux heat wave. Shown are the predictor values the algorithm fashioned in the final iteration that produced grid cells with superb fitness in the metric of surface temperature gain scores. The predictor values for the extreme heat wave are in the left column. The predictor values for faux heat wave 1 are in the right column. Comparisons between the two columns can be used for variable selection in a simple application of feature engineering.151515 This is an example of cherry picking that can invalidate conventional statistical inference. However, the analyses to follow are able sidestep this problem.
@@
Compared to the faux ideal heat wave, the true ideal heat wave is characterized by substantially higher grid cell, gain score elevation (i.e. 623 meters versus 396 meters) and higher grid cell gain score atmospheric pressure at the tropopause (i.e. 11,767 hPa versus 6,729 hPa). The ideal faux heat wave is characterized by higher grid cell, gain score atmospheric temperatures closer to the ground and lower grid cell, gain score atmospheric temperatures at high altitudes. The role of the atmospheric water vapor mass mixing ratios is inconsistent. There is now perhaps a bit better sense of which predictors are related most strongly to a extreme heat waves compared to a faux heat waves.
8 Forecasting Extreme Waves Using a Binary Outcome
@@
We have arrived at a pivotal research destination. If the forecasting goal is help inform policy options two weeks (or more) before an extreme heat wave arrives, one must anticipate how forecasting could undertaken in practice. To be useful, the predictors must be measured and readily available well in advance of a possible heat wave. This implies ongoing data collection on a regular basis.161616 There are also second order requirements. For example, the data collection requires continuity in naming conventions, which factors measured, the temporal and spatial scales, the measurement technology used, and data file formats. If any of these change, forecasting skill can be badly damaged. At the very least, abrupt variation in the data over time will likely be introduced. In addition, promising predictors should not be constant over time. For each grid cell, altitude above sea level, for instance, is a constant that will not change whether or not an extreme heat wave subsequently occurs. It has no forecasting capacity.
@@
Consider again briefly the operationalization an extreme heat wave discussed earlier. An extreme heat wave is a discrete event; it is either present or not present. The task becomes forecasting a categorical binary outcome. Moreover, when using earlier temperature gain scores, we were fitting a consequence extreme heat wave precursors and perhaps the impact of a heat dome. We were not explicitly forecasting a discrete, extreme heat wave.
@@
Returning to the design in Figure 2 for the June true extreme heat wave, we drop the pre-test and recode the post-test to equal “1.” For purposes of comparison, we use July, 2021 as the setting for a faux heat wave (i.e, faux heat wave 2). The pre-test is dropped and the post-test is recoded to equal “0.” The same 144 grid cells are used for a new extreme heat wave design and for the new faux heat wave design; the same grid cells are exposed to one climate regime in June and then another climate regime in July. The June data are then “stacked” above (or below) the July data to form a single dataset. We have an explicit crossover design including both an extreme heat wave and a faux heat wave. The classification format is balanced, and the rare event problem seems resolved.
@@
But the rationale goes a bit deeper. When the data are stacked, each study unit once again serves as its own control. There is no longer a need to compute a pre-test score. We use the last 4 days of July for the faux heat wave because the time offset between the actual and fax heat wave is now 30 days. The chance of spillover effects becomes very small. This helps justify the absence of a pre-test. Because the earlier results for Faux Heat Wave 1 and Faux Heat Wave 2 were so similar, the different seasonal conditions between June and July probably do not matter much for our classification analysis. Finally, by replacing a regression random forest with a classification random forest, we set the stage for latter of estimating forecasting uncertainty with conformal prediction sets (i.e., not regions).
@@
Random forests was retrained as a classifier using the new data structure. All of the predictors used previously were candidates for the classification exercise. As a start, however, we selected three predictors from Table 2 that seemed most promising: (1) atmospheric pressure at the tropopause, (2) atmospheric water vapor mass mixing ratio at level 8, and (3) the temperature at level 8. For each, we considered the gaps shown in Table 2 between values in the actual heat wave column and values in the faux heat wave column. For the mixture and temperature variables, other proximate altitudes could have been selected instead because proximate altitudes tend to be substantially correlated. Dependence between the selected three predictors by themselves was not a serious complication.
No causal claims are being made. Our earlier analyses are consistent with the expectation that relatively few predictors can suffice for promising classification accuracy. High classification accuracy can be a precursor to high forecasting accuracy.171717 Classification accuracy is not the same as forecasting accuracy. For classification, the outcomes of interest are known and included in the data. Using a statistical classifier, the goal is to correctly identify each known outcome. For forecasting, the outcomes are unknown. If they were known, there would be no need to forecast them. The goal is to correctly project what the outcomes will be. Although as a statistical matter, high classification accuracy can support high forecasting accuracy, forecasting introduces additional assumptions and uncertainties that must be properly addressed. But high classification accuracy does not necessarily imply credible causal inference.
@@
Our random forests results are encouraging. The overall classification error rate is 5.0%. In this setting, a false positive can be defined as incorrectly classifying the faux heat wave as the true heat wave. The false positive rate is 5.7%. In this setting, a false negative can be defined as incorrectly classifying the true heat wave as the faux heat wave. The false negative rate is about 4.1%. In short, the random forest classifier almost always correctly identifies the true positives and the true negatives. This can enormously benefit conformal prediction inference.
@@
Figure 5 is the variable importance plot from random forests. The metric at the bottom of the figure is the reduction in classification accuracy average over both outcome classes when each variable in turn is randomly shuffled with the other variables fixed at their empirical existing values. For example, classification accuracy declines by nearly 30 percentage points if temperature at altitude 8 (only) is shuffled. In short, Figure 5 shows all three predictors are important for classification accuracy.181818 Average classification accuracy is used because the false negative and false positive rates are very similar. There is at this point no need to focus on either outcome class by itself.
@@
Figure 6 show the partial dependence plots for the three predictors. A partial dependence plot displays the conditional mean of the fitted values for different values of a given predictor, all other predictors held constant at their mean values in the data (Friedman, 2001; Greenwell, 2017). These values necessarily are unchanging as the selected predictor’s values change; they are fixed. In this way, the possible impact of correlated predictors is controlled, but not with conventional covariance adjustments used parametric classification (e.g., logistic regression).191919 Unlike covariance adjustments, no assumptions need be made about how the predictors are related to one another. The vertical axis in Figure 6 is in probability units with black dots the fitted values. The blue line overlaid is a lowess smoother (Cleveland and Devlin, 1988).
@@
The partial dependence plot at the top is for the predictor atmospheric temperature at altitude 8. Judging by the smoothed fitted values, there is an S-Shaped relationship with the probability of a heat wave but the fitted values are actually flat until the temperature is approximately 232 K and after it reaches about 235 K. Sparse data at in the tails of the temperature distribution may be the explanation. Overall, the relationship is positive, and ranges from near 0.0 to near 1.0 with temperature increases two weeks earlier.
@@
The partial dependence plot in the middle is for the atmospheric pressure at the tropopause. Judging by the smoothed fitted values, there is as atmospheric pressure rises nearly a straight line increase in the probability of a heat wave from about a 0.0 probability to almost a 1.0 probability. However, fitted values are flat until atmospheric pressure reaching 11500 hpa, perhaps an impact of sparse data for lower atmospheric pressures.
@@
The bottom partial dependence plot shows the relationship between heat wave probabilities and the atmospheric water vapor mass mixing ratio at altitude 8. Consistent with its weak importance to the fit, the relationship shown by the smoother is seems haphazard. The fitted values appear disorderly. However, the disappointing performance does not seem to undermine the overall excellent classification accuracy
8.1 Forecasting Uncertainty and Conformal Prediction Sets
Given the data available and possible peculiarities of PNW grid cells, classification accuracy is very promising. A next task is to turn classification into forecasting coupled with valid estimate of uncertainty. We use nested conformal prediction sets. For an unlabeled grid cell, conformal prediction sets can produce an extreme heat wave and/or a no extreme heat wave forecast that is provably correct at a predetermined probability. Because the random forest classifier discussed earlier was so accurate, there is good reason to expect that such forecasts will be very accurate as well. To address uncertainty in practical manner, we proceed with nested conformal prediction sets under the popular “split sample” approach (Kuchibhotla and Berk, 2023).
The key assumption is that the dataset used is comprised of exchangeable observations. This means that the order in which the observations are realized does not matter for the joint probability distribution of the data. When forecasts are needed for new cases without labels, the new observations also are realized in an exchangeable manner from the same population as the data on hand.
Our crossover design buttresses the exchangeable assumption. It does not matter whether the data from June were realized in July and data from July were realized in June. Their joint probability distribution would be the same. Recall that because of the 30 day disparity between the June actual heat wave and July faux heat wave, spillover effects are highly unlikely, and there is no other apparent connection in the data between the late heat wave events in June and the late faux heat wave events in July.202020 The issues can be subtle. This means that if for the data on hand the designated months were swapped, the histogram for the full dataset would be unchanged. It does not mean that grid cells measured in different months are exposed to the same climate conditions.
As the split sample approach requires, the data on hand were randomly split into two disjoint samples, one used to train a classifier, and one used to construct conformal prediction sets. Some call the second split a “test” sample. Some call it “calibration” data. One advantage of this approach is that most any classifier can be used, even methods like stepwise logistic regression, which aggressively capitalize on data snooping that can invalidate convention statistical inference (Berk et al., 2013). Such problems and others are not transported from the training data to the calibration data because the calibration dataset only gets to work with the classifier output, not the processes by which the output was produced. The calibration data distribution provides a baseline used to evaluate hypothetical forecasted outcomes for unlabeled data for which a forecast is sought.212121 Technically, only the calibration data need to be exchangeable.
Nested conformal prediction sets (Gupta et al., 2022) was applied to the crossover data for June and July of 2021. A predetermined coverage probability of .75 was specified. When for a given, unlabeled, grid cell a claim is made that the true future outcome is included in the prediction set, it will be correct at least 75% of the time; the odds in favor of the truth of that grid cell’s forecast are at least 3 to 1.
For illustrative purposes, we used 10 grid cells from the training data as if their true outcome class was unknown. For each grid cell, there was a single member in the prediction set, which represents ideal precision. Based on the outcome classes in the calibration data, the forecasts for all 10 cases were correct. This is no surprise because the random forests classifier identified the outcome classes so well. In retrospect, a more demanding coverage probability than .75 could have used (e.g. .90) because the random forests classifier was so accurate.

### v3

@@ 7.1 Constructing Algorithmic Survival Functions
A random forest was trained using 500 regression trees for the June extreme heat wave. The same gain score data were used for training as in the earlier gain score comparisons in Section 5.1. The trained and averaged ensemble of regression trees serves as a survival function for a genetic algorithm’s construction of an ideal, synthetic population of grid cells experiencing a true heat wave.
@@
In the same manner, a random forest was trained as using 500 regression trees on June faux heat wave 1. The trained and averaged ensemble of regression trees serves as a different survival function for a genetic algorithm’s construction of an ideal, synthetic population of grid cells experiencing a faux heat wave.
7.2 An Application of a Genetic Algorithm To Construct An Idea Type True Heat Wave and and Ideal Type Faux Heat Wave
@@
The two survival functions were used as input to the procedure GA in R written by Luca Scrucca (2013). The standard defaults for numeric predictors worked well. For both survival functions, 5000 iterations were specified leading to 5000 synthetic populations of 100 grid cells each having increasing fitness on the average over iterations. As a first approximation, one can think of a genetic algorithm as jittering predictor values, computing new fitted values from the survival function using the jittered predictors, saving the subset of the grid cells with the larger gain scores, and repeating this process many times until there is a population that includes only very fit grid cells (i.e., having large gain scores). In the process, predictor dependence can be reduced substantially because of the randomness the genetic algorithm introduces to predictors.111111 The GA software in R has several tuning parameters with good default values.(1) For each synthetic population of grid cells, the predictors from the five grid cells with the highest computed surface temperatures gain scores were not altered. (2) The cross-over probability for each possible pair of predictors was set at 0.8. A cross-over refers to a pair of “parent” grid cells combining their predictor values to form an “offspring” grid cell. (3) Because all of the predictors were treated as numeric in this application, the predictor values for an offspring grid cell were computed as the average of the corresponding predictors from a random pair of parent grid cells. (4) The mutation probability for each grid cell’s predictor was set at 0.1. (5) The mutation method used replaced the value of the existing numeric variable with a value within a specified range drawn at random from a uniform distribution. The size of a synthetic population is not a tuning parameter but 100 grid cells seemed a good compromise between an ability to examine the output visually in detail and having a sufficient number of observations to arrive at sufficiently stable parameter estimates.
7.2.1 Genetic Algorithm Results
@@
The grid cell mean gain score for the final synthetic population produced by the actual heat wave survival function was 7.26 degrees K. The grid cell mean gain score for the final synthetic population produced by the faux heat wave 1 survival function was 5.76 degrees K. Because of differences in the two survival functions, the genetic algorithm did not equate the two. This is another vehicle through which to reflect on whether the physics for a true heat wave may differ somewhat from the physics of a faux heat wave.
@@
Predictor True Heat Wave Faux Heat Wave 1
@@
Table 2 provides more evidence about the ways in which an ideal true heat wave can differ from an ideal faux heat wave. Shown are the predictor values the algorithm fashioned for the final population that produced grid cells with superb fitness in the metric of surface temperature gain scores. The predictor values for the extreme heat wave are in the left column. The predictor values for Faux Heat Wave 1 are in the right column. Comparisons between the two columns can be used for variable selection in a simple application of feature engineering.121212 This is an example of cherry picking that can invalidate conventional statistical inference. One is engaged, even if indirectly, in model selection (Berk et al., 2013). However, analyses to follow are able sidestep this problem.
@@
Compared to the ideal faux heat wave, the ideal true heat wave is characterized by (1) substantially higher grid cell elevation (i.e. 623 meters versus 396 meters), (2) a higher tropopause (i.e. 11,767 m versus 6,729 m), (3) higher grid cell atmospheric temperatures at lower altitudes, and (4) lower grid cell atmospheric temperatures at higher altitudes. The role of the atmospheric water vapor mass mixing ratios is inconsistent. There is now perhaps a bit better sense of which predictors are related most strongly to an extreme heat waves compared to a faux heat waves.
A minor complication is that grid level elevation and location over land or water are both fixed. They cannot be directly related to the extreme heat wave. The genetic algorithm was applied a second time, but the two fixed variables were not included. There were no material changes in the conclusions compared to those from Table 2, and the mean gain scores for the final populations were effectively the same as well.
8 Extreme Heat Wave Classification Using a Binary Outcome
@@
With the long term goal accurate forecasting of extreme heat waves two weeks (or more) in advance, several practical hurdles exist. The predictors must be measured and readily available well in advance of a possible heat wave. This implies data collection on a regular basis. There are also several second order requirements. For example, the data collection requires continuity in naming conventions, variables measured, the temporal and spatial scales, the measurement technology used, and data file formats. If any of these change, forecasting skill can be badly damaged. At the very least, abrupt nuisance variation in the data over time will likely be introduced.
In addition, a pre-test is no longer feasible. Any pre-test immediately before a possible heat wave makes forecasting unworkable. When a forecast is needed two weeks in advance, the pre-test values would not yet have materialized. Data curation and associated research designs must adjust.
@@
Moreover, a numeric outcome such as surface temperature should be revisited. An extreme heat wave is a discrete event; it is either present or not present. Therefore, the task becomes forecasting a binary outcome. When using earlier temperature gain scores, we were fitting a consequence extreme heat wave precursors and perhaps the impact of a heat dome. We were not explicitly forecasting a discrete, extreme heat wave. Regression should be replaced by classification.
Finally, as an intermediate step in the direction of real forecasting, it should be possible for predictors located substantially earlier to identify accurately real heat waves compared to faux heat waves. But accurate classification does not guarantee accurate forecasting; for classification the outcome is known whereas when forecasting, the outcome is not known. (If it were known, there would be no need for a forecast.) New statistical complications can arise.
8.1 Research Design Revisions
@@
Returning to the design in Figure 2 for the June true extreme heat wave, we drop the pre-test. We also recode the post-test to equal “1.” For purposes of comparison, we use July, 2021 as the setting for a faux heat wave (i.e, Faux Heat Wave 2). The pre-test is dropped and the post-test is recoded to equal “0.” The same 144 grid cells are used for a new extreme heat wave design configuration and for the new faux heat wave design configuration; the same grid cells are exposed to one climate regime in June and then another climate regime in July. The June data are then “stacked” above (or below) the July data to form a single dataset. The classification format is balanced, and the rare event problem seems resolved.
But the rationale goes a bit deeper. When the data are stacked, each study unit no longer serves as its own control. What remains, however, is that the composition of the study units does not change. The same set of grid cells are exposed precursors of a true heat wave and precursors of the faux hear waves. The grid cells serve as the their own collective controls. The confounding would be far more challenging if one set of grid cells were exposed to true heat wave precursors and another set of grid cells were exposed to faux heat wave precursors.
@@
We use the last 4 days of July for the faux heat wave because the time offset between the actual and fax heat wave is 30 days. The chance of carry-over effects becomes very small. This helps justify the absence of a pre-test. Because the earlier results for Faux Heat Wave 1 and Faux Heat Wave 2 were so similar, the different seasonal conditions between June and July probably do not matter much for our classification analysis.
8.2 Classification Using Random Forests
The earlier comparisons between predictor distributions for a subsequent true heat wave or a subsequent faux heat wave coupled with the genetric algorithm output promise excellent classification results. We considered two forms of widely used and successful supervised learning classifiers: random forests (Breiman, 2001a) and stochastic gradient boosting (Friedman, 2001). Both can be seen as variants of nonparametric regression, are very easy to train, and often the default tuning parameter settings work well. A complication is that there is strong dependence between many of the AIRS predictors. For example, the measured temperatures at different altitudes are often correlated at values above .95.
Random forests handles strongly correlated predictors relatively well because it is comprised of a large ensemble of regression or classification trees with the partitioning for each tree undertaken using a small random subsets of predictors. Sampling predictors (without replacement) for each split can dramatically reduce the impact of high multicollinearity.131313 Random forests usually randomly samples cases when each tree is grown and predictors at each split in a tree. Random sampling of cases can be done with or without replacement. Random sampling predictors and cases is designed to increase the independence between trees. Greater independence facilitates the averaging over all trees in the forest. We recently learned that gbm3 in R now allows sampling predictors too, and along its R predecessor gbm also allows random sampling of cases. In R, the choice between random forests (i.e., randomForests) gradient boosting now depends largely on details that are peripheral to this discussion. We settled on the implementation of random forests in R, ported by Andy Law and Matthew Wiener (Liaw and Weiner, 2002) from the original Fortran by Breiman and Cutler (Breiman, 2001a), which includes several important auxiliary algorithms that can dramatically increase algorithmic transparency.141414 We briefly considered using deep neural networks through the keras3 package in R authored and maintained by Tomasz Kalinowski. Unlike random forests algorithm, neither the predictors nor the observations are randomly sampled. The vastly different units for some of the predictors caused problems for the back propagation. Dependencies between the predictors also caused computational difficulties. Once both problems were addressed, tuning was able to proceed. For all practical purposes, the fitting performance was about as good as random forests’. But, some of the auxiliary procedures available for random forests, which shine some light into the algorithmic black box (Molnar, 2022), were not readily available. For example, the widely applicable partial dependence plots procedure pdp, authored and maintained by Brandon Greenwell, apparently requires that a custom prediction function be coded when used on neural networks output.
@@ 8.2 Classification Using Random Forests
Random forests was trained as a classifier using the new data structure. All of the predictors used previously were candidates for the classification exercise. As a start, however, we selected three predictors from Table 2 that were dynamic and seemed most promising: (1) height at the tropopause, (2) atmospheric water vapor mass mixing ratio at level 8, and (3) the temperature at level 8. For each, we considered the gaps shown in Table 2 between values in the synthetic heat wave column and values in the synthetic faux heat wave column. For the mixture and temperature variables, other proximate altitudes could have been selected because proximate altitudes tend to be substantially correlated. Dependence between the three selected predictors by themselves was not a terribly serious complication but still risks some unstable results. We emphasize again that an algorithm is not a model and even if it were, that model would be badly misspecified. There is also the matter of cherry picking the predictors.
@@
The random forests results are rather strong. Overall classification error rate is approximately 5.0%. In this setting, a false positive can be defined as incorrectly classifying the faux heat wave as the true heat wave. The false positive rate is 5.7%. In this setting, a false negative can be defined as incorrectly classifying the true heat wave as the faux heat wave. The false negative rate is about 4.1%. In short, the random forest classifier almost always correctly identifies the true positives and the true negatives (i.e., very high sensitivity and specificity respectively).
@@
Figure 5 is the variable importance plot from random forests. The metric at the bottom of the figure is the reduction in classification accuracy averaged over both outcome classes when each variable in turn is randomly shuffled with the other variables fixed at their existing values. For example, classification accuracy declines by nearly 30 percentage points if temperature at altitude 8 (only) is shuffled. In short, Figure 5 shows that all three predictors are important for classification accuracy.151515 Average classification accuracy over both outcome classes is used because the numbers of false negatives and false positive are very similar. There is at this point no need to distinguish between the two outcome classes.
@@
Figure 6 show the partial dependence plots for the three predictors. On each vertical axis is the fitted conditional probability of the extreme heat wave for different values of a given predictor, all other predictors fixed at their mean values in the data (Friedman, 2001; Greenwell, 2017).161616 The probabilities range from 0.0 to 1.0 with fitted values needing much of that range. Placing all three plots in the same figure (in the interests of space) attenuates visually the strength of the associations summarized by the smoother. The possible impact of correlated predictors used by the classifier is controlled, but not with conventional covariance adjustments used parametric classification (e.g., logistic regression). The blue line overlaid is a lowess smoother (Cleveland and Devlin, 1988).
@@
The partial dependence plot at the top is for the predictor atmospheric temperature at altitude 8. Judging by the smoothed fitted values, there is an S-shaped relationship with the probability of a heat wave, but the fitted values are actually flat until the temperature is approximately 232 K and flat again after the temperature reaches about 235 K. Sparse data in the tails of the temperature distribution may be the explanation. Overall, the relationship is positive. Fitted probabilities range from near 0.0 to near 1.0 with temperature increases from two weeks earlier.
@@
The partial dependence plot in the middle is the height of the tropopause in meters. Judging by the smoothed fitted values, there is as the height increases nearly a straight line increase in the probability of a heat wave from about a 0.0 probability to almost a 1.0 probability. However, fitted values are flat until atmospheric pressure reaches 11500 meters, perhaps an impact of sparse data for lower atmospheric pressures.
@@
The bottom partial dependence plot shows the relationship between heat wave probabilities and the atmospheric water vapor mass mixing ratio at altitude 8. Consistent with its weaker importance to the fit, the relationship shown by the smoother seems haphazard. The fitted values appear disorderly. However, the disappointing performance does not seem to undermine the overall classification accuracy when all three variables are used at once.
In summary, the classification exercise produced superb accuracy. Clearly, the predictors work even with a 2 week lag. The variable importance plot and the partial dependence plots help to make sense of the accuracy. However, our extensive exploratory examination of the data places a thumb on accuracy scale. The empirical results are valid on their face, but in other settings, absent the such a rich preliminary examination of the data, classification accuracy could be substantialy lower.
8.3 Forecasting Uncertainty
As a statistical matter, high classification accuracy can support high forecasting accuracy. However, in addition to the complications just described, we picked the true heat wave dates before the classification analysis began. Our results are conditional on first knowing the dates of the true heat wave and the faux heatwaves and then assigning the appropriate outcome classes homogeneously to all grid cells. That alone makes forecasting problematic. In particular, one cannot properly take the next step into valid forecasting.
One should ordinarily expect proper forecasts coupled with estimates of uncertainty. A good option for discrete outcomes is conformal prediction sets that produce legitimate forecasts and valid uncertainty estimates even in finite samples. The key assumption is exchangeability (Gupta et al., 2022). Under the popular split samples approach, the data are randomly separated into two disjoint samples: (1) a training dataset that a classifier can use for fitting, and (2) a calibration dataset used construct prediction sets. A wide variety of classifiers can be used be they algorithms or models. There is no requirement that either contains all of the relevant predictors or employs the substantively appropriate functional forms. Still, for unlabeled cases needing a forecast, a conformal prediction set will include a forecast at a predetermined coverage probability properly interpreted as the probability that the forecast is correct. No assumptions need be made about the joint probability distribution of the dataset or additional properties of the calibration data.
But, one must be able to make a credible case that at least the calibration data are exchangeable. This is difficult requirement for the PNW data. Because we selected the outcomes in advance, the realized data are not IID or exchangeable, and the calibration data would not be exchangeable either. In addition, the algorithm specification incorporated by the classifier is the product of various kinds of prior analyses raising post-model-selection inferential challenges (Berk et al., 2013; Kuchibhotla et al., 2022). But it is still useful to consider whether the results of the exploratory data analysis are replicable.